### v1

# Voltage-tuned anomalous-metal to metal transition in hybrid Josephson junction arrays

Abstract
We report voltage-tuned phase transitions in arrays of hybrid semiconductor-superconductor islands arranged in a square lattice. A double-layer electrostatic gate geometry enables independent tuning of inter-island coupling and proximity-induced superconductivity. This design enables access to the superconductor-insulator, superconductor-metal, and metal-insulator transitions in a single device, revealing critical points and emergent intermediate phases. We find that the superconductor-insulator transition is interrupted by an anomalous metallic phase with saturating low-temperature resistivity. Across gate voltages, this regime extends over three orders of magnitude in resistivity and can be continuously tuned into the conventional metallic phase. The signature of the anomalous metallic phase is suppressed by magnetic frustration.
Two-dimensional systems were long thought to transition directly from a superconducting to an insulating phase with increasing effective disorder [1, 2, 3, 4, 5]. However, numerous experiments have since reported an intermediate metallic phase characterized by finite, saturating resistivity as temperature, $T$, approaches zero [6, 7]. This anomalous metallic (AM) phase has been observed in various systems, including thin films [8, 9, 10, 11, 12], oxides [13, 14], crystalline superconductors [15], patterned islands on graphene [16], perforated high-$T_{\mathrm{C}}$ superconductors [17], and Josephson junction arrays [18, 19, 20, 21].
Several mechanisms have been proposed to explain the AM behavior, including order parameter fluctuations [22, 23, 24], emergence of a Bose metal [25, 26, 27], coupling to a dissipative environment [28, 29], formation of a composite Fermi liquid [30], residual heating due to insufficient electronic filtering [31], and free vortex motion [32, 33]. As a result, recent studies have shifted focus from the specifics of the superconductor-insulator transition (SIT) to the broader question of how a metallic state can persist at low $T$ when superconductivity is suppressed by external perturbations [34, 35, 36]. This includes understanding how the anomalous metal relates to a conventional metal, and how its extent depends on the degree of disorder.
Here, we investigate voltage-tuned phase transitions in two-dimensional semiconductor-superconductor hybrid Josephson junction arrays. The system features dual electrostatic gates that enable independent control of the coupling between neighboring islands and the coupling between the islands and the surrounding semiconductor. This geometry enables experimental access to superconducting, insulating, and metallic states within a single device. By decreasing the inter-island coupling, we observe an SIT interrupted by an AM phase, with low-temperature saturation resistivity that can be tuned across more than three orders of magnitude around the superconducting resistance quantum, $h/4e^{2}$. When tuning the proximity-induced superconductivity independently, the AM phase continuously evolves into a conventional metallic phase. Finally, we find that frustrating the array by applying a perpendicular magnetic field introduces the $T$ dependence in the resistivity, suppressing the signatures of the AM phase. We attribute this behavior to thermally activated vortex motion.
The measured hybrid Josephson junction arrays were fabricated on an InAs two-dimensional electron gas (2DEG) heterostructure proximitized by epitaxial Al, similar to that described in Refs. [37, 38]. The Al layer was lithographically patterned into a $255\times 31$ array of $1\,\mu$m square islands with 400 nm spacing. After covering the structure with a thin HfOx dielectric, a 100 nm wide frame gate was metallized between the Al islands to tune their coupling; see Fig. 1(a). A second HfOx layer and a global Ti/Au top gate were added to control the uncovered regions around the islands. A schematic cross-section of the complete device stack is shown in Fig. 1(b). An optical image of the final device, overlaid with the measurement setup is shown in Fig. 1(c). Two additional devices, with different island sizes and spacings, showed similar results. In the main text, we focus on data from a representative device. Supporting data from the other two devices are presented in the Supplementary Material [39]. Measurements were performed using standard lock-in techniques in a dilution refrigerator with a three-axis vector magnet and base temperature of 15 mK. Further details on the growth, sample fabrication, and measurements are provided in the Supplemental Material [39].
Figure 1: (a) Colorized scanning electron micrograph of a reference device, taken before depositing the global top gate. The square Al islands (gray) are patterned on top of the semiconducting heterostructure (green-gray) and are separated by a frame gate (yellow). (b) Schematic cross-section of the device illustrating dual gate geometry. The lower frame gate tunes the central part of the junctions, while the global top gate tunes the 2DEG surrounding the islands. (c) Optical micrograph of the measured Hall bar device showing measurement setup.
Figure 2: (a) Map of resistivity, $\rho$, measured for the main hybrid array as a function of top-gate, $V_{\mathrm{TG}}$, and frame-gate, $V_{\mathrm{FG}}$, voltages, showing the superconducting, insulating, and metallic states. (b) Traces of $\rho$ as a function of $V_{\mathrm{FG}}$ for different temperatures, $T$, and fixed $V_{\mathrm{TG}}=-0.5$ V showing a crossing point around $h/4e^{2}$ indicating superconductor-insulator transition. (c) Similar to (b) but as a function of $V_{\mathrm{TG}}$ at fixed $V_{\mathrm{FG}}=-0.3$ V with a crossing point suggesting superconductor-metal transition.
We begin by characterizing the resistivity, $\rho$, of the hybrid array as a function of frame-gate, $V_{\mathrm{FG}}$, and top-gate, $V_{\mathrm{TG}}$, voltages at base temperature; see Fig. 2(a). For $V_{\mathrm{FG}}=0$ and $V_{\mathrm{TG}}=0$, the array is in a global superconducting state characterized by vanishing $\rho$. This reflects coherent coupling between proximitized regions beneath neighboring islands via the Josephson effect.
Both gate voltages reduce the Josephson coupling between Al islands, but their effect on the global state of the device differs considerably. When decreasing $V_{\mathrm{FG}}$, we observe a transition into a highly resistive state as carriers in the junctions between islands are depleted, eventually disconnecting the islands. The temperature evolution of $\rho(V_{\mathrm{FG}})$ at fixed $V_{\mathrm{TG}}=-0.5$ V reveals a crossing point of all traces around $V_{\mathrm{FG}}^{*}=-1.17$ V, where $\rho$ is roughly $h/4e^{2}$ and shows negligible $T$ dependence; see Fig. 2(b). Below this point, $\rho$ increases with decreasing $T$, whereas above it, $\rho$ decreases. This behavior is consistent with an SIT, comparable to previous observations in similar hybrid arrays [19]. In contrast, a negative $V_{\mathrm{TG}}$ depletes the regions around the islands, weakening the proximity effect while maintaining conduction through the 2DEG under the frame gate. At sufficiently low $V_{\mathrm{TG}}$, the islands become effectively isolated and the array develops a finite $\rho$. The corresponding $\rho(V_{\mathrm{TG}})$ traces also show a $T$-independent crossing point around $V_{\mathrm{TG}}^{*}=-2.2$ V, though with weaker temperature dependence on the resistive side, confirming a superconductor-metal transition; see Fig. 2(c). Depending on the value of $V_{\mathrm{TG}}$, sweeping $V_{\mathrm{FG}}$ can induce either a superconductor-insulator or a metal-insulator transition.
Figure 3: (a) Resistivity, $\rho$, as a function of temperature, $T$, measured for different frame-gate voltage, $V_{\mathrm{FG}}$, values ranging from $-1.060$ to $-1.225$ V, at fixed $V_{\mathrm{TG}}=0$. The dashed black lines are linear fits of $\ln(\rho)$ versus $\ln(T)$ up to $70$ mK. (b) Low-$T$ resistivity slope, $d\ln(\rho)/d\ln(T)$, taken from the fits in (a), plotted against the average resistivity, $\bar{\rho}$, in the same $T$ range. Error bars are the standard uncertainties of the fitting.
To better understand the nature of the voltage-tuned SIT, we investigate the temperature dependence of $\rho$ at different values of $V_{\mathrm{FG}}$ while keeping $V_{\mathrm{TG}}=0$; see Fig. 3(a). For all $V_{\mathrm{FG}}$ values, a small step in $\rho$ is visible around $T=1.8$ K, signifying the superconducting transition of the thin Al islands [40]. As $T$ is decreased, traces at $V_{\mathrm{FG}}\gtrsim-1.12$ V show a sharp drop in $\rho$ that eventually falls below the experimental sensitivity $(\lesssim 10^{-4}\,h/4e^{2})$, confirming a global superconducting state. For slightly more negative $V_{\mathrm{FG}}\approx-1.14$ V, $\rho$ still drops initially but then saturates below $T\approx 100$ mK. The saturation resistivity increases continuously with decreasing $V_{\mathrm{FG}}$, spanning over three orders of magnitude (from roughly $10^{-3}$ to $10~{}h/4e^{2}$). For $V_{\mathrm{FG}}\lesssim-1.20$ V, $\rho$ starts to increase with decreasing $T$, indicating a crossover to the insulating state of the device.
To characterize the saturating regime at low temperature, we analyze the low-$T$ slope of $\ln(\rho)$ versus $\ln(T)$ for different $V_{\mathrm{FG}}$. For each trace, we perform a linear fit to the data below 70 mK, yielding $d\ln(\rho)/d\ln(T)$. A positive (negative) slope indicates a superconducting (insulating) state. The intermediate regime, with a suppressed slope but finite $\rho$, can be identified as the anomalous metal [6]. This can be illustrated by plotting $d\ln(\rho)/d\ln(T)$ as a function of average resistivity, $\bar{\rho}$, in the fitted $T$ range; see Fig. 3(b). The slope is positive for $\bar{\rho}\lesssim 10^{-2}~{}h/4e^{2}$ and negative for $\bar{\rho}\gtrsim 10^{2}~{}h/4e^{2}$, consistent with superconducting and insulating states, respectively, as $T\to 0$. For $\bar{\rho}$ between $10^{-2}$ and $10~{}h/4e^{2}$, $d\ln(\rho)/d\ln(T)$ is roughly zero, indicating that the SIT is interrupted by the AM phase, spanning three orders of magnitude of $\bar{\rho}$.
We note that $\rho$ at $V_{\mathrm{FG}}=-1.20$ V remains roughly $h/4e^{2}$ throughout the measured $T$ range [Fig. 3(a)]. For less negative $V_{\mathrm{FG}}$, $\rho$ drops below its normal-state value—in some cases by several orders of magnitude—before saturating at low $T$. For slightly more negative $V_{\mathrm{FG}}$, the first couple of traces still saturate, but at values higher than their high-$T$ resistivity. This behavior resembles a smooth crossover from a failed superconductor to a failed insulator within the anomalous metallic phase and may be attributed to quantum fluctuations [41, 34, 42].
The AM phase is observed as $V_{\mathrm{FG}}$ reduces the carrier density in the junctions between the islands, driving the SIT, while the regions around the islands remain populated ($V_{\mathrm{TG}}=0$). However, as mentioned earlier, depleting those regions can tune the system into a conventional metallic phase. With this in mind, we investigate how the system evolves between these two metallic regimes. Specifically, we focus on the narrow region that connects the intermediate-$\rho$ regime associated with the anomalous metal at less negative $V_{\mathrm{TG}}$ to the broader metallic regime emerging at more negative $V_{\mathrm{TG}}$; see Fig. 2(a). This connection is seen more clearly in a zoomed-in map of $\bar{\rho}(V_{\mathrm{FG}},V_{\mathrm{TG}})$, averaged over $T=15$ to 55 mK range taken in 2.5 mK steps; see Fig. 4(a).
Figure 4: (a) Average resistivity, $\bar{\rho}$, measured as a function of top-gate, $V_{\mathrm{TG}}$, and frame-gate, $V_{\mathrm{FG}}$, voltages, showing an enlargment of the region where the superconducting, insulating, and metallic states meet. The plot was constructed by averaging a set of $\rho$ maps over the temperature range from $T=15$ to 55 mK, taken every $2.5$ mK. (b) Low-$T$ $\rho$ slope, $d\ln(\rho)/d\ln(T)$, obtained by linear fitting of $\ln(\rho)$ versus $\ln(T)$ for the same data as in (a)
To further probe how the anomalous metal evolves into a conventional metal, we use the same data set to construct a map of the slope $d\ln(\rho)/d\ln(T)$; see Fig. 4(b). At the top-right corner of the map, global superconductivity yields vanishing resistance across the entire $T$ window, resulting in zero slope. More negative gate voltages gradually suppress Josephson coupling, leading to finite resistivity with a positive slope [red regions in Fig. 4(b)]. The left side of the map is dominated by noise, indicating that the islands become disconnected from the source and drain. However, just before complete decoupling, the array exhibits insulating behavior with a negative slope [blue regions in Fig. 4(b)]. The slope remains near zero between the superconducting and insulating boundaries, with no disruption along the narrow region as $V_{\mathrm{TG}}$ decreases. This continuity demonstrates a smooth evolution from anomalous to conventional metallic states, despite the distinct mechanisms driving the associated transitions.
Figure 5: (a) Resistivity, $\rho$, measured as a function of temperature, $T$ and magnetic frustration, $f$, shows peaks of the superconducting transition temperature around commensurate frustration values, mostly pronounced around $f=0$ and $1/2$. (b) Traces of $\rho$ as a function $T$, taken for three representative frustration values, $f=0$, irrational ($\approx 0.16$), and $f=1/2$, at fixed $V_{\mathrm{FG}}=-1.13$ V. The normal-state resistivity matches for all the traces, but nonzero $f$ show higher $\rho$ at low $T$ with finite slope. (c) Similar to (b) but at $V_{\mathrm{FG}}$ chosen such that $\rho$ at the base-$T$ match for all frustrations.
All data discussed so far were taken at magnetic field carefully tuned to zero (for more details, see Supplemental Material [39]). We now turn to investigate the effect of an out-of-plane magnetic field, $B$, which introduces a superconducting phase gradient between neighboring islands [43]. To this end, we map the resistivity as a function of $T$ and magnetic frustration, $f=\Phi/\Phi_{0}$, where $\Phi=BA$ is the magnetic flux through a unit cell of area $A$ and $\Phi_{0}=h/2e$ is the superconducting flux quantum. We find that the critical temperature, $T_{\mathrm{C}}$, is maximal at $f=0$ and decays quickly as the array is frustrated; see Fig. 5(a). Additional peaks in $T_{\mathrm{C}}$ appear at several other frustration values, most prominently at $f=1/2$, corresponding to commensurate vortex configurations.
To probe how frustration affects the AM phase, we compare $\rho(T)$ traces at $f=0$, irrational ($\approx 0.16$), and $1/2$ under two conditions. First, we select traces with matching normal-state resistivity; see Fig. 5(b). At $f=0$, $\rho$ saturates at low $T$, characteristic of the anomalous metal. In contrast, the traces with nonzero $f$ show a finite positive slope down to the base temperature, lying well above the $f=0$ trace. Second, we keep the same $f=0$ trace and select traces with matching base-temperature resistivity; see Fig. 5(c). Here, the low-$T$ slope at nonzero $f$ remains finite, and their normal-state resistivity is lower than at $f=0$.
In both cases, the emergence of a small but finite positive slope at nonzero $f$ indicates that magnetic frustration destabilizes the low-temperature saturation of the AM phase. We attribute this behavior to thermally activated vortex motion, which introduces dissipation even before the onset of the Berezinskii–Kosterlitz–Thouless transition [21]. The observation of the finite slope confirms that the system remains thermally sensitive down to the base temperature, supporting an intrinsic AM phase at $f=0$ rather than an artifact of residual heating.
In summary, we have investigated voltage-controlled phase transitions in Josephson junction arrays patterned on an InAs/Al heterostructure. Using a dual-gate geometry, we find a robust anomalous metallic phase interrupting the superconductor-insulator transition, which can be smoothly tuned into a conventional metallic state with fully decoupled superconducting islands. At zero magnetic frustration, the anomalous metallic phase, characterized by vanishing low-temperature dependence, persists across more than three orders of magnitude in resistivity. A finite frustration suppresses the resistivity saturation, indicating thermally activated vortex motion and confirming thermal sensitivity down to the base temperature.
Acknowledgments—We thank S. Upadhyay for assistance with the device fabrication and L. Banszerus, M. R. Lykkegaard, A. Padley, and T. Røhling for discussions. We acknowledge support from research grants (Projects No. 43951, No. 53097, and No. 50334) from VILLUM FONDEN, the Danish National Research Foundation, the European Research Council (Grant Agreement No. 856526), and the MEYS grant (Project No. LM202351).